a boy is 10 years old his father is 37 years old.in haw many years time will the father be twice as old as his son
Answer by ikleyn(52798)

.
x+37 = 2(x+10).

Solve for x.

x = 37 - 2*10 = 37 - 20 = 17.

Answer. In 17 years.